# OpenGait: A Comprehensive Benchmark Study for Gait Recognition towards Better Practicality

### v1

@@ V-A Pipeline
As shown in Fig. 4 (a), the backbone transforms each input frame into a 3D feature map with the height, width, and channel dimensions. Then, a temporal pooling (TP) module aggregates the obtained feature map sequence by performing the maximization along the temporal dimension, outputting a global understanding. Next, the obtained feature map is horizontally divided into several parts, and each part is pooled into a feature vector, according to the popular horizontal pooling (HP) [47]. As a result, we get several feature vectors and use separate fully connected layers (FCs) to map them into the metric space, i.e., generating $f$. Since these part vectors will be processed independently, the following formulation loosely treats $f$ as a single feature vector for brevity.
@@
Given a mini-batch $\mathbb{B}$, we can get the collection of positive and negative pairs over this batch, namely $\mathbb{S}^{+}$ and $\mathbb{S}^{-}$. For each sample pair $(\overset{\to}{f},\overset{\leftarrow}{f})$, we utilize the Euclidean norm to measure their distance, i.e., $d=\left\|\overset{\to}{f}-\overset{\leftarrow}{f}\right\|$. Then the triplet loss is employed to drive the identity learning process:
$L_{triplet}=\frac{1}{\mathcal{R}(\mathbb{S}^{+})}\frac{1}{\mathcal{R}(\mathbb{S}^{-})}\sum_{i\in\mathbb{S}^{+}}\sum_{j\in\mathbb{S}^{-}}\left\lfloor d_{i}-d_{j}+m\right\rfloor_{+}$ (1)
@@
where $\mathcal{R}(\cdot)$ measures the collection size and $m$ presents the loss margin. Furthermore, we use BNNecks [55] to optimize the identity distribution by estimating the probability $p_{\hat{y}}$ for class $\hat{y}$. Let the ground-truth label be $y$ and class number be Y, then the cross-entropy loss can be formulated as:
$p_{y}$ $=p_{\hat{y}}\quad\text{if}\quad\hat{y}=y\quad\text{else}\quad 1-p_{\hat{y}}$ (2)
$L_{ce}$ $=-\frac{1}{\mathcal{R}(\mathbb{B})}\sum^{\mathbb{B}}\frac{1}{\text{Y}}\sum^{\text{Y}}_{y}\text{log}(p_{y})$
In the end, the overall loss is $L=L_{triplet}+L_{ce}$.
V-B Appearance-based DeepGaitV2
Gait data is typically presented as sparse input, such the silhouette image with a size of $64\times 44$, which might create the impression that gait recognition is an ‘easy’ task that shallow networks can handle, as researchers used to do. In the context of developing baseline models, we agree that the gait recognition defined by constrained gait datasets, such as the CASIA-B [10], OU-MVLP [11], and the latest SUSTech1K [21] and CCPG [20] datasets, can indeed be considered as relatively simple cases that shallow gait models can deal with. This is because they merely contain limited covariates manipulated by researchers, whereas gait models may circumvent human-made challenges by focusing on evidently unchanged cues rather than subtle walking patterns [20]. Therefore, we assume that deep gait models may easily over-fit on in-the-lab datasets. To validate this hypothesis, we build DeepGaitV2 series.
Figure 5: The DeepGaitV2 series meets the over-fitting cases on (a) CASIA-B and (b) OU-MVLP, with the network depth increasing. The loss number presents the count of triplets that cause non-zero loss in the training batch, directly reflecting the network’s convergence state.
@@ V-B Appearance-based DeepGaitV2
Down-sampling is performed by Stage 2 and 3 with a stride of 2. $T$ and $C$ respectively denote the input sequence length and arbitrary channel number.
@@
TABLE VII: The Rank-1 Accuracy of DeepGaitV2 with Various Depths.
Method Depth In-the-lab In-the-wild
@@
GaitSet [7] $<$10 87.1 87.1 36.7 48.4
GaitPart [15] $<$10 88.5 88.7 28.2 47.6
GaitGL [16] $<$10 91.9 92.0 29.7 47.3
GaitBase [17] 10 89.4 90.8 60.1 60.1
DeepGaitV2 (Shallow to Deep) 10 89.6 92.0 66.8 69.3
14 85.5$\downarrow$ 92.0 70.8$\uparrow$ 75.7$\uparrow$
22 75.3$\downarrow$ 91.9$\downarrow$ 72.8$\uparrow$ 79.4$\uparrow$
30 65.7$\downarrow$ 91.2$\downarrow$ 71.7$\downarrow$ 79.5$\uparrow$
Here the rank-1 accuracy on CASIA-B is the average over its three conditions.
With deepening the backbone, as illustrated in Table VII, there exists a noticeable distinction in performance trends between the constrained and real-world evaluations. Specifically, applications on Gait3D [9] and GREW [8] exhibit substantial benefits from properly deep networks, while contrasting outcomes are observed on CASIA-B [10] and OU-MVLP [11]. The analysis shown in Fig. 5 (a) and (b) uncover the over-fitting issues on CASIA-B and OU-MVLP, i.e., the 22-layer DeepGaitV2 converges better but performs worse.
Moreover, we find similar over-fitting phenomena on the latest constrained gait datasets, namely SUSTech1K [21] and CCPG [20]. Despite the introduction of new human-controlled challenges such as poor illumination, complex occlusion, and clothing changes, as shown in Fig. 5 (c) and (d), the 22-layer DeepGaitV2 continues to exhibit the over-fitting behavior.
Combining the above observations, this paper advocates for the adoption of deep models as the fundamental architecture for analyzing in-the-wild gait datasets in future research endeavors. According to Table VII, we propose utilizing the 10-layer and 22-layer DeepGaitV2 models as baseline models for constrained and real-world evaluations, respectively. Beyond deepening DeepGaitV2, we also develop a light version for better accuracy-speed balance. Specifically, we employ pseudo 3D block [57] shown in Fig. 4 (b) to reconstruct DeepGaitV2 shown in Table VI. This slight change can save about 59% training weights and 57% computation costs777 # param: 27.5 vs. 11.1 MB, and FLOPs: 6.8 vs. 2.9 GFLOPs per frame. Here only consider the backbone, and the same below unless otherwise stated. with remaining competitive performances on both Gait3D [9] (+1.6%) and GREW [8] (-1.7%). In the following, we use the pseudo-3D DeepGaitV2 as the default unless otherwise stated.
Figure 6: The network architectures of DeepGaitV2 v.s. SkeletonGait v.s. SkeletonGait++. The ‘head’ part is ignored for brevity.
@@ V-C Model-based SkeletonGait
Previous skeleton-based methods [23, 30, 27] typically model the coordinates of body joints as non-grid graphs using graph neural networks. In this benchmark study, we introduce a novel skeleton-based gait representation called the skeleton map, drawing inspirations from related works [58, 59, 60]. As illustrated in Fig. 4 (c), our skeleton map represents the coordinates of joints as a heatmap, i.e., transforming the skeleton into a silhouette-like image without exact body shapes. This manner offers three significant advantages:
  • Aligning the skeleton and silhouette data formats enables a seamless extension of DeepGaitV2 to skeleton-based methodology, eliminating the need for additional efforts to develop graph-specific gait models. This alignment ensures consistency in the network architecture between silhouette-based and skeleton-based gait recognition, making further comparisons intuitive and fair.
  • The skeleton map, resembling a silhouette without body shapes, allows for an intuitive controlled experiment to observe the distinct roles of body structure vs. shape features used in describing gait patterns for recognition.
  • Serving as an imagery input, the skeleton map can be seamlessly integrated into image-based gait models, particularly at the bottom convolution stages, facilitating the further exploration of multi-modal baseline models in this benchmark study.
As a result, SkeletonGait is developed by replacing the input of DeepGaitV2 from the silhouette to the skeleton map, as shown in Fig. 6 (b). The only architectural modification involves adjusting the input channel, as the silhouette is single-channel while the skeleton map is double-channel. Specifically, we generate the skeleton map through the following steps.
Given the coordinates of human joints $(x_{k},y_{k},c_{k})$, where $(x_{k},y_{k})$ and $c_{k}$ respectively present the location and confidence score of the $k$-th joint with $k\in\{1,...,K\}$ ($K$ denotes the number of body joints). Firstly, considering the absolute coordinates of joints relative to the original image contain much gait-unrelated information like the walking trajectory and filming distance, we introduce the pre-treatments of center- and scale-normalization to align raw coordinates:
@@
Typically, the height of the human body is expected to exceed its width. As a result, the normalized coordinates of human joints, as defined in Eq. (3), should fall within the range of $H\times H$. But in practice, the pose estimator is imperfect and may produce some outlier joints outside the $H\times H$ scope. To address these out-of-range cases, the resolution of the skeleton map, denoted as $R$, should be larger than $H$, ensuring coverage of all the coordinates. In our experiments, let $R$ be $2H$ is enough for all the employed datasets.
As illustrated in Fig. 7 (a), the skeleton map is initialized as a blank image with a size of $R\times R$. Then we draw it based on the normalized coordinates of human joints. Inspired by [58], we generate the joint map J by composing $K$ Gaussian maps, where each Gaussian map is centered at a specific joint position and contributes to all the $R\times R$ pixels:
@@
Similarly, we can also create a limb map L:
@@
Next, the skeleton map is obtained by stacking J and L and thus has a size of $2\times R\times R$. Notably, for the convenience of visualization, we repeat the last channel of all the skeleton maps shown in this paper to display the visual three-channel images with the size of $3\times R\times R$.
Figure 8: More examples of the skeleton coordinates v.s. silhouette images v.s. skeleton maps.
As shown in Fig. 7 (b), we employ subject-centered cropping to remove the blank regions, thus reducing the redundancy in skeleton maps. In practice, the vertical range is determined by the minimum and maximum heights of pixels which possess non-zero values. Meanwhile, the horizontal cropping range spans from $\frac{R-H}{2}$ to $\frac{R+H}{2}$. In this way, we remove extraneous areas outside the desired gait region, ensuring a more concise and compact skeleton map. Lastly, to align with the input size required by downstream gait models, the cropped skeleton maps are resized to $2\times 64\times 64$ and further cropped by the widely-used double-side cutting strategy.
As a result, Fig. 8 exhibits some examples of the used skeleton maps with varying $\sigma$. As we can see, a smaller $\sigma$ produces a visually thinner skeleton map, whereas excessively large $\sigma$ may lead to visual ambiguity.
@@
  • Discriminability. Preceding methods tend to directly resize the obtained images of varying sizes into a predetermined fixed size, inevitably resulting in the loss of body ratio information. Conversely, the scale-normalization and subject-centered cropping techniques outlined in this paper ensure that the skeleton map preserves the authenticity of the length and ratio of human limbs.
@@ V-D Multi-modal SkeletonGait++
To integrate the superiority of silhouette and skeleton map, as shown in Fig. 6 (c), SkeletonGait++ provides a fusion-based two-stream architecture involving the silhouette and skeleton branch. These two branches respectively share the same network architectures with DeepGaitV2 and SkeletonGait at early stages, such as Conv 0 and Stage 1. Then, a fusion module is responsible for aggregating these two feature sequences in a frame-by-frame manner. This benchmark study considers three kinds of fusion mechanisms:
  • Add Fusion. The feature maps from the silhouette and skeleton branch are combined using an element-wise addition operation, as demonstrated in Fig. 6 (d).
@@
  • Attention Fusion. The feature maps from the silhouette and skeleton branch are first concatenated along the channel dimension, and then transformed by a small network to form a cross-branch understanding. Here the small network is composed of a squeezing $1\times 1$, a plain $3\times 3$, and an expansion $1\times 1$ convolution layer. As shown in Fig. 6 (e), a softmax layer is next employed to assign element-wise attention scores respectively for the silhouette and skeleton branch. Lastly, an element-wise weighted-sum operation is used to generate the output.
Next, the Stage 3 and 4 possess the same network architectures as the SkeletonGait. Moreover, we also consider the fusion location. Fig. 6 (c) exhibits the low-level fusion case. Another high-level fusion model aggregates the features before Stage 4, with additional Stage 2 and 3 respectively being inserted into the silhouette and skeleton branch.
@@ VI More Experiments
For a convincing elaboration, the key statistics of employed datasets have been presented in Table I, and some experiments have been briefly described in the previous section along with exploring DeepGaitV2. In this section, we provide the implementation details for completeness, and then exhibit the performance comparison and ablation study.
TABLE VIII: Recognition results on the OUMVLP [11], Gait3D [9], and GREW [8] datasets.
Category Method Venue Testing Datasets
OU-MVLP [11] Gait3D [9] GREW [8]
rank-1 rank-1 rank-5 mAP mINP rank-1 rank-5 rank-10 rank-20
Silhouette- based GaitSet [7] AAAI’19 87.1 36.7 58.3 30.0 17.3 46.3 63.6 70.3 -
GaitPart [15] CVPR’20 88.5 28.2 47.6 21.6 12.4 44.0 60.7 67.3 -
GaitGL [16] ICCV’21 89.7 29.7 48.5 22.3 13.6 47.3 -
DANet [61] CVPR’23 90.7 48.0 69.7 - - -
GaitBase [17] CVPR’23 90.8 60.1 - 64.6 -
GaitSSB [62] T-PAMI’23 91.8 63.6 - 61.7 -
QAGait [63] AAAI’24 - 67.0 81.5 56.5 - 59.1 74.0 79.2 83.2
DeepGaitV2 Ours 91.9 74.4 88.0 65.8 39.2 77.7 88.9 91.8 93.0
Skeleton- based GaitGraph2 [30] CVPRW’22 62.1 11.1 - 33.5 -
Gait-TR [64] ES’23 56.2 6.6 - 54.5 -
GPGait [29] ICCV’23 60.5 22.5 - 53.6 -
SkeletonGait Ours 67.4 38.1 56.7 28.9 16.1 77.4 87.9 91.0 93.2
Multi- modal SMPLGait [9] CVPR’22 - 46.3 64.5 37.2 22.2 -
GaitRef [65] IJCB’23 90.2 49.0 49.3 40.7 25.3 53.0 67.9 73.0 77.5
ParsingGait [25] MM’23 - 76.2 - 68.2 - -
HybridGait [66] AAAI’24 - 53.3 72.0 43.3 26.7 -
SkeletonGait++ Ours - 77.6 89.4 70.3 42.6 85.8 92.6 94.3 95.5
The highest results are in bold for silhouette-based methods and underlined for skeleton-based methods. Regarding multi-modal methods, the highest results are in both bold and underlined. The same annotation is applied in the following tables.
TABLE IX: Some Implementation details.
@@
The batch size $(8,16)$ indicates $8$ subjects and $16$ sequences per subject.
Implementation Details. Table IX displays the main hyper-parameters of our experiments. Unless otherwise specified, a) The silhouettes are aligned by the normalization strategy used in [11] and resized to $64\times 44$. The spatial augmentation strategy outlined by [17] is adopted. b) Different datasets often employ various skeleton data formats, such as COCO 18 for OU-MVLP and BODY 25 for CCPG. To enhance flexibility, our implementation standardized these various formats to COCO 17 uniformly. c) At the test phase, the entire gait sequence is fed into the model directly. As for the training stage, the data sampler collects a fixed-length segment of 30 frames as input. d) The margin $m$ in Eq. (1) is set to 0.2. For the cross-entropy loss shown in Eq. (2), a soft classification trick [67] is employed to encourage the model to be less confident on the training set. e) The SGD optimizer with an initial learning rate of 0.1 and weight decay of 0.0005 is utilized. f) For DeepGaitV2, the $C$ controlling the channel number in Table VI is set to 64 as default. For SkeletonGait, the $\sigma$ controlling the variance in Eq. (4) and (5) is set to 8.0 as default. For SkeletonGait++, the fusion strategy is instantiated as the attention fusion at the low-level stage as depicted in Fig. 6 (c) and (f). g) All the experiments strictly follow the official evaluation protocols. h) All the implementations are based on our OpenGait. All the codes are publicly available at https://github.com/ShiqiYu/OpenGait.
VI-A Comparison Around DeepGaitV2

### v2

@@ V-A The Pipeline
As shown in Fig. 4 (a), the backbone transforms each input frame into a 3D feature map with the height, width, and channel dimensions. Then, a temporal pooling (TP) module aggregates the obtained feature map sequence by max pooling along the temporal dimension, it outputs a global understanding of the data. Next, the obtained feature map is horizontally divided into several parts, and each part is pooled into a feature vector, according to the popular horizontal pooling (HP) in [54]. As a result, we can get several feature vectors and use separate fully connected layers (FCs) to map them into the metric space, i.e., generating $f$. Since these vectors will be processed independently for each part, our formulation loosely treats $f$ as a single feature vector for brevity.
@@
Given a mini-batch $\mathbb{B}$, we can get a collection of positive and negative pairs over this batch, namely $\mathbb{S}^{+}$ and $\mathbb{S}^{-}$. For each sample pair $(f^{+}_{i},f^{-}_{i})$ where $i\in\mathbb{S}^{+}\cup\mathbb{S}^{-}$, we utilize the Euclidean norm to measure their distance, i.e., $d_{i}=\left\|f^{+}_{i}-f^{-}_{i}\right\|$. Then the triplet loss is employed to drive identity learning:
$L_{triplet}=\sum_{i\in\mathbb{S}^{+}}\sum_{j\in\mathbb{S}^{-}}\left\lfloor d_{i}-d_{j}+m\right\rfloor_{+}$ (1)
where $\left\lfloor\cdot\right\rfloor_{+}$ denotes the ReLU function [62] and $m$ is the loss margin [63] with a default value of 0.2. In practice, $L_{triplet}$ is further reduced by the number of non-zero terms over $\left\lfloor\cdot\right\rfloor_{+}$.
@@
Next, we use BNNecks [56] to optimize the identity distribution. For each $i$-th sample within batch $\mathbb{B}$, we let $p^{i}_{\hat{y}}$ present its probability belonging to class $\hat{y}\in\{1,2,...,\text{Y}\}$, where Y is the class number. Given the ground-truth label $y^{i}$, the cross-entropy loss can be formulated as:
$L_{ce}$ $=-\frac{1}{\mathcal{R}(\mathbb{B})}\sum_{i\in\mathbb{B}}\frac{1}{\text{Y}}\sum_{\hat{y}\in\{1,2,...,\text{Y}\}}\text{log}(p^{i}_{y})$ (2)
$p_{y}^{i}$ $=p_{\hat{y}}^{i}\quad\text{if}\quad\hat{y}=y^{i}\quad\text{otherwise}\quad 1-p_{\hat{y}}^{i}$
where $\mathcal{R}(\cdot)$ measures the sample number within batch $\mathbb{B}$. In practice, a soft classification trick [64] is further employed to encourage the model to be less confident on the training set.
@@
Note: Down-sampling is performed by Stage 2 and 3 with a stride of 2. $T$ and $C$ respectively denote the sequence length and arbitrary channel number.
V-B Appearance-based DeepGaitV2
Gait data is usually in a simple format, such as the silhouette with a size of $64\times 44$. It creates the impression that gait recognition is an easy task that shallow networks can handle. In the context of developing baseline models, we agree that the gait recognition simulated by constrained gait datasets, such as the CASIA-B [10], OU-MVLP [11], and the latest SUSTech1K [21] and CCPG [20] datasets, can be considered as relatively simple tasks that shallow gait models can deal with. The reason is that they contain limited covariates manipulated in data collection. The models can circumvent human-made challenges by focusing on evidently unchanged cues rather than subtle walking patterns [20]. Therefore, we assume that deep gait models easily overfit on in-the-lab datasets. To validate this hypothesis, we build the DeepGaitV2 series.
Figure 5: The DeepGaitV2 series meets the over-fitting cases on (a) CASIA-B and (b) OU-MVLP, with the network depth increasing. The loss number presents the count of triplets that cause a non-zero loss in the training batch, directly reflecting the network’s convergence state.
@@ V-B Appearance-based DeepGaitV2
TABLE VII: The Rank-1 Accuracies by DeepGaitV2 with Various Depths.
Method Depth Param. GFLOPs In-the-lab In-the-wild
@@
GaitSet [7] $<$10 2.3 12.9 87.1 87.1 36.7 48.4
GaitPart [15] $<$10 3.8 7.9 88.5 88.7 28.2 47.6
GaitGL [16] $<$10 7.0 58.5 91.9 92.0 29.7 47.3
GaitBase [17] 10 4.9 35.5 89.4 90.8 60.1 60.1
DeepGaitV2 (Shallow to Deep) 10 14.2 91.5 89.6 92.0 66.8 69.3
14 18.6 129.0 85.5$\downarrow$ 92.0 70.8$\uparrow$ 75.7$\uparrow$
22 27.5 203.7 75.3$\downarrow$ 91.9$\downarrow$ 72.8$\uparrow$ 79.4$\uparrow$
30 41.6 278.5 65.7$\downarrow$ 91.2$\downarrow$ 71.7$\downarrow$ 79.5$\uparrow$
Note: the rank-1 accuracy on CASIA-B is the average over its three conditions. Param. refers to model size (M).
With deepening the backbone, as illustrated in Table VII, there exists a noticeable distinction in performance trends between in-the-lab and in-the-wild evaluations. Specifically, applications on Gait3D [9] and GREW [8] exhibit substantial benefits from properly deep networks, while opposite outcomes are observed on CASIA-B [10] and OU-MVLP [11]. The analysis shown in Fig. 5 (a) and (b) uncover the over-fitting issues on CASIA-B and OU-MVLP, i.e., the 22-layer DeepGaitV2 converges better but performs worse.
Moreover, we find similar over-fitting phenomena on the latest constrained gait datasets, SUSTech1K [21] and CCPG [20]. Despite the introduction of new human-controlled challenges such as poor illumination, complex occlusion, and clothing changes, as shown in Fig. 5 (c) and (d), the 22-layer DeepGaitV2 continues to exhibit the over-fitting behavior.
Combining the above observations, this paper advocates for the adoption of deep models as the fundamental architecture for analyzing in-the-wild gait datasets in future research endeavors. According to Table VII, we utilize the 10-layer and 22-layer DeepGaitV2 models as baseline models for in-the-lab and in-the-wild evaluations, respectively. Beyond deepening DeepGaitV2, we also develop a light version for better accuracy-speed balance. Specifically, we employ pseudo 3D block [58] shown in Fig. 4 (b) to reconstruct DeepGaitV2 shown in Table VI. This slight change can save about 59% training weights and 57% computation costs888 # param: 27.5 vs. 11.1 MB, and FLOPs: 6.8 vs. 2.9 GFLOPs per frame. Here we only consider the backbone (the same below unless otherwise stated). with remaining competitive performances on both Gait3D [9] (+1.6%) and GREW [8] (-1.7%). In the following, we use the pseudo-3D DeepGaitV2 as the default unless explicitly stated.
Figure 6: The network architectures of (a) DeepGaitV2, (b) SkeletonGait and (c) SkeletonGait++. The head part is ignored for brevity. The different fusions in SkeletonGait++ are (d) add fusion, (e) cat fusion and (f) attention fusion.
@@ V-C Model-based SkeletonGait
Previous skeleton-based methods typically model the coordinates of body joints as non-grid graphs using graph neural networks [28, 35, 32]. In this benchmark study, we introduce a novel skeleton-based gait representation called the skeleton map, drawing inspirations from related works in [65, 66, 67]. As illustrated in Fig. 4 (c), our skeleton map represents the coordinates of joints as a heatmap, i.e., transforming the skeleton into a silhouette-like image without exact body shapes. This manner offers three significant advantages:
  • Aligning the skeleton and silhouette data formats can enable a seamless extension of DeepGaitV2 to skeleton-based methodology. It eliminates the need for additional efforts to develop graph-specific gait models. This alignment ensures consistency in the network architecture between silhouette-based and skeleton-based gait recognition. So it makes further comparisons intuitive and fair.
  • The skeleton map, resembling a silhouette without body shapes, allows for an intuitive controlled experiment to observe the distinct roles of body structure vs. shape features in gait recognition.
  • Serving as an imagery input, the skeleton map can be seamlessly integrated into image-based gait models, particularly at the bottom convolution stages. It can facilitate the further exploration of multi-modal baseline models.
As a result, SkeletonGait is developed by replacing the input of DeepGaitV2 from silhouettes to skeleton maps, as shown in Fig. 6 (b). The only difference is the number of input channels, as the silhouettes are single-channel while the skeleton maps are double-channel. Specifically, the skeleton map can be generated by the following steps.
Given the coordinates of human joints $(x_{k},y_{k},c_{k})$, where $(x_{k},y_{k})$ and $c_{k}$ respectively present the location and confidence score of the $k$-th joint with $k\in\{1,...,K\}$ ($K$ denotes the number of body joints). Firstly, considering the absolute coordinates of joints relative to the original image contain much gait-unrelated information like the walking trajectory and filming distance, we introduce a center-normalization and a scale-normalization to align the coordinates:
@@
Typically, the height of the human body is expected to exceed its width. As a result, the normalized coordinates of human joints, as defined in Eq. (3), should fall within the range of $H\times H$. But in practice, the pose estimator is imperfect and may produce some outlier joints outside the $H\times H$ scope. To address these out-of-range cases, the resolution of the skeleton map, denoted as $R$, should be larger than $H$. Then all the coordinates will be inside the map. In our experiments, we set $R=2H$ for all datasets.
As illustrated in Fig. 7 (a), the skeleton map is initialized as a blank image with a size of $R\times R$. Then it is drawn based on the normalized coordinates of human joints. Inspired by [65], the joint map J is generated by composing $K$ Gaussian maps, where each Gaussian map is centered at a specific joint position and contributes to all the $R\times R$ pixels:
@@
Similarly, a limb map L can also be created:
@@
Next, a skeleton map can be obtained by stacking J and L, and thus has a size of $2\times R\times R$. Notably, for the convenience of visualization, we repeat the last channel of all the skeleton maps shown in this paper to display the visual three-channel images with the size of $3\times R\times R$.
Figure 8: Skeleton coordinates vs. silhouette images vs. skeleton maps.
As shown in Fig. 7 (b), subject-centered cropping is employed to remove the blank regions. It can reduce the redundancy in skeleton maps. In our experiments, the vertical range is determined by the minimum and maximum heights of pixels that possess non-zero values. Meanwhile, the horizontal cropping range spans from $\frac{R-H}{2}$ to $\frac{R+H}{2}$. In this way, extraneous areas outside the desired gait region can be removed. It ensures a concise and compact skeleton map. Lastly, to align with the input size required by downstream gait models, the cropped skeleton maps are resized to $2\times 64\times 64$ and further cropped by the widely-used double-side cutting strategy.
Fig. 8 presents some examples of the used skeleton maps with varying $\sigma$. It can be found that a smaller $\sigma$ produces a visually thinner skeleton map, whereas excessively large $\sigma$ may lead to visual ambiguity.
@@
  • Discriminability. Directly resizing the obtained images with varying sizes into a predetermined fixed size, which is used in some previous methods, will inevitably result in the loss of body ratio information. Conversely, the scale-normalization and subject-centered cropping introduced in this paper ensure that the skeleton map preserves the authenticity of the length and ratio of human limbs.
@@ V-D Multi-modal SkeletonGait++
To integrate the superiority of silhouettes and skeleton maps, as shown in Fig. 6 (c), SkeletonGait++ provides a fusion-based two-stream architecture with a silhouette branch and a skeleton one. These two branches respectively share the same network architectures with DeepGaitV2 and SkeletonGait at their early stages, such as Conv 0 and Stage 1. Then, a fusion module is to aggregate the two feature sequences from the two branches in a frame-by-frame manner. Three kinds of fusion mechanisms are considered and evaluated in the experiments.
  • Add Fusion. The feature maps from the silhouette and skeleton branches are combined using an element-wise addition operation, as demonstrated in Fig. 6 (d).
@@
  • Attention Fusion. The feature maps from the silhouette and skeleton branch are first concatenated along the channel dimension, and then transformed by a small network to form a cross-branch understanding. Here the small network is composed of a squeezing $1\times 1$, a plain $3\times 3$, and an expansion $1\times 1$ convolution layer. As shown in Fig. 6 (f), a softmax layer is employed to assign element-wise attention scores respectively for the silhouette and skeleton branches. Lastly, an element-wise weighted-sum operation is used to generate the output.
In this work, SkeletonGait++ considers two fusion locations. The first is a low-level counterpart as illustrated in Fig. 6 (c), where the fusion module works before Stage 2. Another high-level fusion aggregates the features before Stage 4, with additional Stage 2 and 3 respectively being inserted into the silhouette and skeleton branches.
@@ VI More Experiments
For a convincing elaboration, the key statistics of employed datasets in the experiments are presented in Table I. The experimental designs have been briefly described in the previous section along with the description of DeepGaitV2. In this section, the implementation details are provided for completeness, and then the comparison and ablation studies are presented.
TABLE VIII: Some Implementation details.
@@
Note: batch size $(8,16)$ denotes $8$ subjects and $16$ samples per subject.
Implementation Details. Table VIII lists the main hyper-parameters of our experiments. Unless otherwise specified, the settings are as follows.
  • The silhouettes are aligned by the normalization strategy in [11] and resized to $64\times 44$. The spatial augmentation strategy outlined in [17] is adopted.
  • Different datasets may have different skeleton formats, such as COCO 18 for OU-MVLP and BODY 25 for CCPG. To enhance flexibility, in our experiments all formats are standardized to COCO 17.
  • At the test phase, a whole gait sequence is fed into the model directly whatever how many frames it has. But for training the data sampler collects a fixed-length clip of 30 frames as the input.
  • The SGD optimizers are with an initial learning rate of 0.1 and a weight decay of 0.0005. The learning rate will be reduced by 0.1 at each milestone in Table VIII.
  • For DeepGaitV2, the channel $C$ in Table VI is set to a default value of 64. For SkeletonGait, the variance $\sigma$ in Eq. (4) and (5) is set to a default value of 8.0. The fusion strategy in SkeletonGait++ is instantiated as the low-level attention fusion as depicted in Fig. 6 (c) and (f).
  • All the recognition metrics mentioned below are estimated according to the evaluation protocol provided by the dataset creators. For instance, the Rank-$k$ (R-$k$) accuracy represents the identification rate compared with the top $k$ ranking samples in the gallery set. Specifically for the Gait3D dataset, the mean Average Precision (mAP) and mean Inverse Negative Penalty (mINP) [68], which consider the recall of multiple instances and hard samples, have also been included in the comparison.
  • All the implementations are by OpenGait available at https://github.com/ShiqiYu/OpenGait.
TABLE IX: Recognition results on the OU-MVLP [11], Gait3D [9], and GREW [8] datasets.
Category Method Year Testing Datasets
OU-MVLP [11] Gait3D [9] GREW [8]
Rank-1 Rank-1 Rank-5 mAP mINP Rank-1 Rank-5 Rank-10 Rank-20
Silhouette- based GaitSet [7] 2019 87.1 36.7 58.3 30.0 17.3 46.3 63.6 70.3 -
GaitPart [15] 2020 88.5 28.2 47.6 21.6 12.4 44.0 60.7 67.3 -
GaitGL [16] 2021 89.7 29.7 48.5 22.3 13.6 47.3 -
DANet [69] 2023 90.7 48.0 69.7 - - -
GaitBase [17] 2023 90.8 64.6 - 60.1 -
GaitSSB [44] 2023 91.8 63.6 - 61.7 -
QAGait [70] 2024 - 67.0 81.5 56.5 - 59.1 74.0 79.2 83.2
DeepGaitV2 Ours 91.9 74.4 88.0 65.8 39.2 77.7 88.9 91.8 93.0
Skeleton- based GaitGraph2 [35] 2022 62.1 11.1 - 33.5 -
Gait-TR [71] 2023 56.2 6.6 - 54.5 -
GPGait [34] 2023 60.5 22.5 - 53.6 -
SkeletonGait Ours 67.4 38.1 56.7 28.9 16.1 77.4 87.9 91.0 93.2
Multi-modal SMPLGait [9] 2022 - 46.3 64.5 37.2 22.2 -
GaitRef [72] 2023 90.2 49.0 49.3 40.7 25.3 53.0 67.9 73.0 77.5
HybridGait [73] 2024 - 53.3 72.0 43.3 26.7 -
SkeletonGait++ Ours - 77.6 89.4 70.3 42.6 85.8 92.6 94.3 95.5
Note: The highest results are in bold for silhouette-based methods and underlined for skeleton-based methods. Regarding multi-modal methods, the highest results are in both bold and underlined. The same annotations are applied in the following tables.
VI-A Analysis by the Appearance-based DeepGaitV2